# Automated Coral Spawn Monitoring for Reef Restoration: The Coral Spawn and Larvae Imaging Camera System (CSLICS)

### v1

@@ Abstract
Coral aquaculture for reef restoration requires accurate and continuous spawn counting for resource distribution and larval health monitoring, but current methods are labor-intensive and represent a critical bottleneck in the coral production pipeline. We propose the Coral Spawn and Larvae Imaging Camera System (CSLICS), which uses low cost modular cameras and object detectors trained using human-in-the-loop labeling approaches for automated spawn counting in larval rearing tanks. This paper details the system engineering, dataset collection, and computer vision techniques to detect, classify and count coral spawn. Experimental results from mass spawning events demonstrate an F1 score of 82.4% for surface spawn detection at different embryogenesis stages, 65.3% F1 score for sub-surface spawn detection, and a saving of 5,720 hours of labor per spawning event compared to manual sampling methods at the same frequency. Comparison of manual counts with CSLICS monitoring during a mass coral spawning event on the Great Barrier Reef demonstrates CSLICS’ accurate measurement of fertilization success and sub-surface spawn counts. These findings enhance the coral aquaculture process and enable upscaling of coral reef restoration efforts to address climate change threats facing ecosystems like the Great Barrier Reef.
@@ I Introduction
The Great Barrier Reef (GBR) is one of the world’s most important and diverse ecosystems, serving as a habitat for over 25% of all marine species and as a nursery for over 75% [1]. Coral reefs also provide social, economic, and cultural services with an estimated value of over 1 trillion USD globally [1, 2]. However, coral reefs around the world are severely threatened by climate change [3]. Prolonged exposure to abnormally high ocean temperatures causes mass coral bleaching and death, resulting in a projected 70-90% decrease in live coral on reefs by 2050 if no action is taken [4, 5]. If coral reefs are to survive these accelerating pressures, scalable and effective restoration approaches are urgently required.
Coral aquaculture has emerged as one such approach, enabling large-scale sexual reproduction of corals in controlled facilities [6]. Unlike asexual fragmentation-based methods, sexual reproduction can produce orders of magnitude more corals while promoting genetic diversity and adaptive capacity to environmental change [6, 7, 8]. The long-term success of coral aquaculture, however, depends critically on monitoring and managing the millions of early-stage coral spawn and larvae in culture tanks. Automated and continuous monitoring is therefore essential to ensure that this otherwise scalable method delivers viable recruits at the scale required for meaningful reef restoration.
@@
The coral aquaculture process involves broodstock selection, spawn collection, fertilization and larval rearing, settlement, and reseeding to reefs [9, 10]. A critical bottleneck lies in larval rearing, where millions of fragile eggs and embryos must be monitored during their early life stages [7]. Quantitative measures, such as fertilization success, are vital for predicting culture viability, yet current monitoring relies on manual sampling and visual counts under a microscope—a process which takes up to 20 minutes per sample. With large facilities projected to operate over 60 rearing tanks by 2030, manual monitoring is infeasible: operators are often limited to sampling only a subset of tanks once per day. This data sparsity, combined with the fact that cultures can deteriorate within hours without intervention, makes manual monitoring both labor-intensive and inadequate for large-scale restoration.
@@ II Related Work
Monitoring coral spawn traditionally relies on manual observation and counting, a process that is time-consuming, vulnerable to human error and disruptive to the delicate coral spawn. There have been some recent works exploring the use of technology at different stages of the coral production process [11, 10], and recent advances in deep learning have further expanded the potential for automated monitoring, however existing approaches are not designed specifically for coral spawn monitoring at the early, free-swimming stage – a key bottleneck in coral production. This section first describes the process of manual coral spawn monitoring, reviews advances in automated coral spawn monitoring, and outlines the application of deep learning for counting tasks and annotating imagery.
@@ II-A Manual Coral Spawn Monitoring
In marine science, coral spawn counting is predominantly performed manually due to the complexity of the task, the delicate nature of the developing coral larvae, and the typically small scale of operations [12]. During the critical annual mass-spawning events, consistency in counting is essential, yet manual methods can vary based on the sampling process and the individual performing the counts.
Manual spawn counting in coral aquaculture involves sampling from the tank and then counting from the samples, as illustrated in Fig. 2. Therefore, traditional sampling requires stirring the tank to ensure larvae are uniformly distributed, followed by extraction of six 5 mL samples. Each sample is then manually counted under a stereo microscope, which takes on average 20 minutes per tank sampled. Ideally, sampling would be performed hourly over the seven day larval rearing period, to enable implementation of mitigating strategies during rapid culture declines. However, the manual effort required typically limits this to once daily per tank. The repeated tank stirring can also have a negative impact on the coral spawn, and therefore culture health [12].
@@ II-B Automated Coral Spawn Monitoring
Previous works on automated spawn counting have been limited by cost and functional usability. Randall et al. used large particle flow cytometry [^2] Cytometry is a laboratory technique that measures the number and characteristics of cells. to count coral spawn; however, the required flow cytometer was prohibitively expensive (i.e. $\approx$500,000 USD) and the methods were not easily scalable or integrated with aquaculture facility tanks [12]. Mullen et al. developed an underwater microscope with a microscopic objective in front of an electrically tunable lens for long-term benthic ecosystem studies [13], which was later improved in [14], with the incorporation of focal scans and estimates of photosynthetic efficiency. Shahani et al. used a variable objective lens in their underwater benthic microscope to achieve a higher depth of field imagery via focus stacking [15]. While these solutions [13, 14, 15] demonstrate advances in underwater microscopy, they are not suitable for automated spawn monitoring or designed for integration in large-scale aquaculture facilities.
Several commercially available scientific and industrial imaging products exist for use in aquatic environments (e.g. for monitoring plankton biodiversity) [16, 17, 18]. Recent advances include compact integrated cameras with embedded computer chips (e.g. Luxonis, Arducam, Point Grey). Some of these commercial sensors have been used in the field at macro scales for coral reef conservation projects, including for mapping reefs [19] and distributing coral larvae across reefs [20]. While potentially useful for field applications, the optics used in these systems are unsuitable for coral conservation aquaculture workflows, due to the microscopic scales of coral spawn imaging. Additionally, most commercially-available housings lack the robustness needed to operate in saltwater environments.
@@ II-C Deep Learning for Scientific Counting Applications
Deep learning and computer vision has increased the capabilities of counting-based applications in many scientific monitoring applications, including for cell counting in health applications [21, 22, 23]. Adoption of these technologies has led to ten times faster operations and yields more reliable and consistent results compared to manual counting [21].
In marine science and aquaculture, deep learning has been successfully used for counting plankton eggs [24], shrimp eggs [25], and fish [26, 27], however, it has not yet been used for coral spawn counting.
Semi-automated human-in-the-loop annotation reduces the burden of fully manual labeling by combining model predictions with iterative human validation [28, 29, 30]. This approach has been applied in ecology, including labeling large camera trap datasets and segmenting coral reef imagery [31, 32, 33]. In this work, we adopt active learning with a human-in-the-loop to efficiently annotate coral spawn imagery collected by CSLICS.
@@ III-A Coral Aquaculture with CSLICS
To automate and increase the frequency of coral counting, we propose the following CSLICS process and timings during aquaculture monitoring of early coral life stages for species belonging to the family Acroporidae (see Fig. 3). The process is outlined in the context of using 500 L conical fiberglass larval culture tanks supplied with flow-through ($\approx$125 L/h) filtered seawater (0.04 $\mu m$, typically 26-29°C) [9, 11].
@@ III-C Software Design
When powered on, a camera bring-up service starts the CSLICS system, which configures and triggers the camera via the Robot Operating System (ROS) [34]. Captured images are saved with metadata and then processed by the appropriate surface or sub-surface coral detection model.
We train two YOLOv8 object detectors [35, 36], which were selected for their accessibility, ease-of-use, widely supported framework, extensive documentation, and the variety of model sizes available. The YOLOv8x 640p model was chosen as a suitable balance between inference speed and accuracy, enabling detection on the CSLICS hardware at 2 seconds per frame. We train one model for surface detection of coral spawn, which is capable of detecting coral spawn at different stages of embryogenesis (Fig. 5 and described in Section IV-B). We also train a model for sub-surface detection of the corals. During this phase, it is no longer necessary to differentiate between developmental stages, so this is a single-class coral detector.
We leverage a semi-supervised, human-in-the-loop labeling approach to reduce the annotation effort required and to rapidly label a dataset for training the detection models [37]. We manually annotate a small number of images to train a preliminary object detection model, which is iteratively updated with additional samples obtained by reviewing predictions from the model. Further details on the datasets and labeling approaches are available in Section IV-B.
@@ IV-A Data Collection
Approximately 24 hours after tank stocking, the operators increased the tank aeration to encourage homogeneity of coral density throughout the tank (Fig. 3). At this point, CSLICS were lowered to submerge the front element of the units no more than 1 cm into the water, to minimize interaction with the coral spawn. This marked the beginning of the sub-surface operation. The exact time was recorded by the operators, and CSLICS were then left to continuously capture and process data until the completion of the larval rearing process (6.5 days).
@@ IV-B2 Sub-surface Data
When the corals are dispersed in the water column, the limited depth of field of the optics becomes strongly apparent (e.g. Fig. 5). Corals too close or too far from the camera appear out-of-focus, while those in-focus remain sharp and defined. This is why we perform single-class coral detection for sub-surface imaging. The exact embryogenesis stage is also less critical at this point as the developing corals approach their final developmental stage (planula larvae) with a smooth and oval appearance. Importantly, the sub-surface detector only counts corals that are in focus, and we therefore manually label this data using CVAT [^6] https://www.cvat.ai/, and ensure that only in-focus corals are annotated.
@@ IV-C4 Sub-surface Tank Counts
Figure 6: Fertilization rate over time for two tanks. Left: a successful larval culture with a Root Mean Square Error (RMSE) of 0.101. Right: an unsuccessful larval culture with RMSE 0.131. The figures demonstrate that CSLICS is able to clearly identify fertilization success and provide early warning of deteriorating cultures.
@@ V Results
We evaluate our CSLICS approach for monitoring coral spawn in aquaculture for both the surface and sub-surface operational modes. When in surface mode, we report the surface detections and the fertilization success (Section V-A). For sub-surface operation, we report the sub-surface detections and the total tank counts (Section V-B). For both fertilization success and total spawn counts, we compare against human-derived estimates. Overall, we find that CSLICS effectively monitors rates of fertilization and can clearly differentiate between successful and unsuccessful larvae cultures.
@@ V-A1 Coral Development Stage Detection
We report the performance of our multi-class coral detector in Table I. Our detector achieves greater than 72.8% F1 score for all development classes, and 48.1% for the damaged class. This class suffers in performance due to the relatively low number of examples (839 instances, compared with 1,700-5,300 instances for the other classes) in the training set, as well as the challenging and highly variable visual characteristics of damaged corals. As the damaged class is not included in the fertilization success ratio calculations, we also report the overall F1 score excluding this class, which is 82.4%.
V-A2 Fertilization Success
CSLICS was effectively used to image the surface of the larval rearing tanks and detect corals of different embryonic development stages. From these detections, Eq. 1 was used to determine the fertilization success. Results are shown in Fig. 6, which shows the fertilization rate over time for two tanks, each equipped with CSLICS.
A rolling mean for the fertilization success rate is shown in orange, where the mean is calculated with a window of 20 images (thus averaged over $\approx$3 minutes). This figure demonstrates CSLICS capability to predict a successful fertilization for a tank: the left plot shows the curve for a successful fertilization, as the rate of fertilized to non-fertilized eggs increases steadily over time. The right plot instead shows an unsuccessful larval culture, as indicated by a fertilization rate which remains significantly lower than the successful culture for the duration (Fig. 6). This occurs as unfertilized eggs begin to dissolve and cause a chain reaction throughout the larval culture.
V-B Sub-surface Operation
V-B1 Coral Detection
We report the precision, recall, F1 score and mAP@0.5 of our coral sub-surface detector model in Table II. These results demonstrate that we are able to detect and count coral spawn in the sub-surface setting, with an F1 score of 65.3% and 64.0% mAP@0.5. The sub-surface setting is more challenging than the surface case (as seen in Fig. 5). As the corals reach the settlement stage (Fig. 3), they begin to descend in the tank, and therefore become out-of-focus of the camera. Further, as the camera housing is submerged in the tank, it can become coated with coral larvae residue. This limitation will be addressed in the next iteration of the system design, which will feature a camera imaging the tank through a viewing window on the side of the tank.
@@ V-B1 Coral Detection
YOLOv8 [36] 64.9 65.7 65.3 64.0
V-B2 Sub-surface Total Tank Counts
In sub-surface operation we quantify the spawn counts for the entire tank. The process for estimating total spawn count was described in Section IV-C4. Fig. 7 shows CSLICS sub-surface total tank counts for a culture in December 2023. The rolling mean tank count is shown with a centered window of 40 images, with one standard deviation bounds shown in the lighter colors, and the manual counts are shown as the square markers. To the best of our knowledge, the manual counts were consistent as they were sampled and counted by the same personnel following the same protocol. This figure demonstrates the accuracy of CSLICS tank count estimates.
Figure 7: CSLICS sub-surface tank counts, demonstrating close alignment between CSLICS estimates and manual counts (RMSE = 45,670 corals), sufficient for monitoring purposes.
VI Conclusion
In this work, we proposed an automated coral spawn and larvae imaging camera system to enhance high-density larval rearing in coral aquaculture. We presented the two operational modes of the system (surface and sub-surface), including the system architecture, hardware specifications, software implementation, and the data collection and annotation methods. We trained two models for counting corals of species belonging to the family Acroporidae: one tailored to detecting coral spawn on the surface across different developmental stages, and the other designed for sub-surface spawn detection. The data captured by CSLICS illustrated clear trends between successful and unsuccessful fertilization success ratios to assess the viability of the coral spawn cultures. Additionally, CSLICS provided accurate tank counts for the sub-surface operation, showing comparable trends with manual counts.
CSLICS’ automation and vision-based approach reduced coral disturbance and enabled recording frequent observations into a digital database. This data supports correlation analysis of treatment strategies with larval rearing outcomes. CSLICS enables robust and consistent tracking of coral spawn at different stages of development. Further, CSLICS streamlines coral rearing processes and avoids the 5,720 hours of labor per spawning event that would be required for manual sampling at the same frequency.
The next iteration of the CSLICS system is currently in development, and addresses limitations of this system. The new system will improve detection of damaged cells, and feature side-mounted imaging to reduce the impacts of coral larvae residue build-up on the housing during sub-surface operation.

### v2

@@ Abstract
Coral aquaculture for reef restoration requires accurate and continuous spawn counting for resource distribution and larval health monitoring, but current methods are labor-intensive and represent a critical bottleneck in the coral production pipeline. We propose the Coral Spawn and Larvae Imaging Camera System (CSLICS), which uses low cost modular cameras and object detectors trained using human-in-the-loop labeling approaches for automated spawn counting in larval rearing tanks. This paper details the system engineering, dataset collection, and computer vision techniques to detect, classify and count coral spawn. Experimental results from mass spawning events demonstrate an F1 score of 82.4% for surface spawn detection at different embryogenesis stages, 83% F1 score for sub-surface spawn detection, and a saving of 5,720 hours of labor per spawning event compared to manual sampling methods at the same frequency. Comparison of manual counts with CSLICS monitoring during a mass coral spawning event on the Great Barrier Reef demonstrates CSLICS’ accurate measurement of fertilization success and sub-surface spawn counts. These findings enhance the coral aquaculture process and enable upscaling of reef restoration efforts to address climate change threats facing ecosystems like the Great Barrier Reef.
@@ I Introduction
The Great Barrier Reef (GBR) is one of the world’s most important and diverse ecosystems, serving as a habitat for over 25% of all marine species and as a nursery for over 75% [heron2017impacts]. Coral reefs also provide social, economic, and cultural services with an estimated value of over 1 trillion USD globally [heron2017impacts, costanza2014changes]. However, coral reefs around the world are severely threatened by climate change [hughes2017global]. Prolonged exposure to abnormally high ocean temperatures causes mass coral bleaching and death, resulting in a projected 70-90% decrease in live coral on reefs by 2050 if no action is taken [de201227, Souter2021Status]. If coral reefs are to survive these accelerating pressures, scalable and effective restoration approaches are urgently required.
Coral aquaculture has emerged as one such approach, enabling large-scale sexual reproduction of corals in controlled facilities [randall2020sexual]. Unlike asexual fragmentation-based methods, sexual reproduction can produce orders of magnitude more corals while promoting genetic diversity and adaptive capacity to environmental change [randall2020sexual, banaszak2023applying, dela2020enhancing]. The long-term success of coral aquaculture, however, depends critically on monitoring and managing the millions of early-stage coral spawn and larvae in culture tanks. Automated and continuous monitoring is therefore essential to ensure that this otherwise scalable method delivers viable recruits at the scale required for meaningful reef restoration.
@@
The coral aquaculture process involves broodstock selection, spawn collection, fertilization and larval rearing, settlement, and reseeding to reefs [pollock2017coral, raine2025ai]. A critical bottleneck lies in larval rearing, where millions of fragile eggs and embryos must be monitored during their early life stages [banaszak2023applying]. Quantitative measures, such as fertilization success, are vital for predicting culture viability, yet current monitoring relies on manual sampling and visual counts under a microscope—a process which takes up to 20 minutes per sample. With large facilities projected to operate over 60 rearing tanks by 2030, manual monitoring is infeasible: operators are often limited to sampling only a subset of tanks once per day. This data sparsity, combined with the fact that cultures can deteriorate within hours without intervention, makes manual monitoring both labor-intensive and inadequate for large-scale restoration.
@@ II Related Work
Monitoring coral spawn traditionally relies on manual observation and counting, a process that is time-consuming, vulnerable to human error and disruptive to the delicate coral spawn. There have been some recent works exploring the use of technology at different stages of the coral production process [severati2024autospawner, raine2025ai], and recent advances in deep learning have further expanded the potential for automated monitoring, however existing approaches are not designed specifically for coral spawn monitoring at the early, free-swimming stage – a key bottleneck in coral production. This section first describes the process of manual coral spawn monitoring, reviews advances in automated coral spawn monitoring, and outlines the application of deep learning for counting tasks and annotating imagery.
@@ II-A Manual Coral Spawn Monitoring
In marine science, coral spawn counting is predominantly performed manually due to the complexity of the task, the delicate nature of the developing coral larvae, and the typically small scale of operations [randall2020rapid]. During the critical annual mass-spawning events, consistency in counting is essential, yet manual methods can vary based on the sampling process and the individual performing the counts.
Manual spawn counting in coral aquaculture involves sampling from the tank and then counting from the samples, as illustrated in Fig. 2. Therefore, traditional sampling requires stirring the tank to ensure larvae are uniformly distributed, followed by extraction of six 5 mL samples. Each sample is then manually counted under a stereo microscope, which takes on average 20 minutes per tank sampled. Ideally, sampling would be performed hourly over the seven day larval rearing period, to enable implementation of mitigating strategies during rapid culture declines. However, the manual effort required typically limits this to once daily per tank. The repeated tank stirring can also have a negative impact on the coral spawn, and therefore culture health [randall2020rapid].
@@ II-B Automated Coral Spawn Monitoring
Previous works on automated spawn counting have been limited by cost and functional usability. Randall et al. used large particle flow cytometry [^2] Cytometry is a laboratory technique that measures the number and characteristics of cells. to count coral spawn; however, the required flow cytometer was prohibitively expensive (i.e. $\approx$500,000 USD) and the methods were not easily scalable or integrated with aquaculture facility tanks [randall2020rapid]. Mullen et al. developed an underwater microscope with a microscopic objective in front of an electrically tunable lens for long-term benthic ecosystem studies [mullen2016underwater], which was later improved in [ben2025benthic], with the incorporation of focal scans and estimates of photosynthetic efficiency. Shahani et al. used a variable objective lens in their underwater benthic microscope to achieve a higher depth of field imagery via focus stacking [shahani2021design]. While these solutions [mullen2016underwater, ben2025benthic, shahani2021design] demonstrate advances in underwater microscopy, they are not suitable for automated spawn monitoring or designed for integration in large-scale aquaculture facilities.
Several commercially available scientific and industrial imaging products exist for use in aquatic environments (e.g. for monitoring plankton biodiversity) [cpics, pi10, ISIIS-DPI]. Recent advances include compact integrated cameras with embedded computer chips (e.g. Luxonis, Arducam, Point Grey). Some of these commercial sensors have been used in the field at macro scales for coral reef conservation projects, including for mapping reefs [dunbabin2020uncrewed] and distributing coral larvae across reefs [mou2022reconfigurablerobots]. While potentially useful for field applications, the optics used in these systems are unsuitable for coral conservation aquaculture workflows, due to the microscopic scales of coral spawn imaging. Additionally, most commercially-available housings lack the robustness needed to operate in saltwater environments.
@@ II-C Deep Learning for Scientific Counting Applications
Deep learning and computer vision has increased the capabilities of counting-based applications in many scientific monitoring applications, including for cell counting in health applications [grishagin2015automatic, morelli2021automating, falk2019u]. Adoption of these technologies has led to ten times faster operations and yields more reliable and consistent results compared to manual counting [grishagin2015automatic].
In marine science and aquaculture, deep learning has been successfully used for counting plankton eggs [weldrick2022promising], shrimp eggs [zhang2021shrimp], and fish [ditria2021annotated, connolly2021improved], however, it has not yet been used for coral spawn counting.
Semi-automated human-in-the-loop annotation reduces the burden of fully manual labeling by combining model predictions with iterative human validation [garcia2023ten, dayoub2017episode, desmond2021semi], which has been applied in ecology, including labeling large camera trap datasets and segmenting coral reef imagery [norouzzadeh2021deep, bakana2023digital, raine2024human]. Here, we adopt human-in-the-loop active learning to efficiently annotate coral spawn imagery collected by CSLICS.
@@ III-A Coral Aquaculture with CSLICS
[…]
@@ III-C Software Design
When powered on, a camera bring-up service starts the CSLICS system, which configures and triggers the camera via the Robot Operating System (ROS) [quigley2009ros]. Captured images are saved with metadata and then processed by the appropriate surface or sub-surface coral detection model.
We train two YOLOv8 object detectors [redmon2016you, ultralyticsyolov8], which were selected for their accessibility, ease-of-use, widely supported framework, extensive documentation, and the variety of model sizes available. The YOLOv8x 640p model was chosen as a suitable balance between inference speed and accuracy, enabling detection on the CSLICS hardware at 2 seconds per frame. We train one model for surface detection of coral spawn, which is capable of detecting coral spawn at different stages of embryogenesis (Fig. 5 and Section IV-B). We also train a model for sub-surface detection of the corals, where it is no longer necessary to differentiate between developmental stages, so this is a single-class coral detector.
We leverage a semi-supervised, human-in-the-loop labeling approach to reduce the annotation effort required and to rapidly label a dataset for training the detection models [mosqueira2023human]. We manually annotate a small number of images to train a preliminary object detection model, which is iteratively updated with additional samples obtained by reviewing predictions from the model. Further details on the datasets and labeling approaches are available in Section IV-B.
@@ IV-A Data Collection
Approximately 24 hours after tank stocking, the operators increased the tank aeration to encourage homogeneity of coral density throughout the tank (Fig. 3). At this point, CSLICS were lowered to submerge the units’ front element no more than 1 cm into the water, to minimize interaction with the coral spawn, marking the beginning of the sub-surface operation. The exact time was recorded by the operators, and CSLICS were then left to continuously capture and process data until the completion of the larval rearing process (6.5 days).
@@ IV-B2 Sub-surface Data
When the corals are dispersed in the water column, the limited depth of field of the optics becomes strongly apparent (e.g. Fig. 5). Corals too close or too far from the camera appear out-of-focus, while those in-focus remain sharp and defined. This is why we perform single-class coral detection for sub-surface imaging. The exact embryogenesis stage is also less critical at this point as the developing corals approach their final developmental stage (planula larvae) with a smooth and oval appearance. Importantly, the sub-surface detector only counts corals that are in focus, and we therefore manually label this data using CVAT (https://www.cvat.ai/), and ensure that only in-focus corals are annotated.
@@ IV-C4 Sub-surface Tank Counts
Figure 6: Fertilization rate over time for two tanks. Left: a successful larval culture with a Root Mean Square Error (RMSE) of 0.101. Right: an unsuccessful larval culture with RMSE 0.131. CSLICS is able to clearly identify fertilization success and provide early warning of deteriorating cultures.
@@ V Results
We evaluate our CSLICS approach for monitoring coral spawn in aquaculture for both the surface and sub-surface operational modes. When in surface mode, we report the surface detections and the fertilization success (Section V-A). For sub-surface operation, we report the sub-surface detections and the total tank counts (Section LABEL:subsec:results-subsurface). For both fertilization success and total spawn counts, we compare against human-derived estimates. Overall, we find that CSLICS effectively monitors rates of fertilization and can clearly differentiate between successful and unsuccessful larvae cultures.
@@ V-A1 Coral Development Stage Detection
We report the performance of our multi-class coral detector in Table IV-A. Our detector achieves greater than 72.8% F1 score for all development classes, and 48.1% for the damaged class. This class suffers in performance due to the relatively low number of examples (839 instances, compared with 1,700-5,300 instances for the other classes) in the training set, as well as the challenging and highly variable visual characteristics of damaged corals. As the damaged class is not included in the fertilization success ratio calculations, we also report the overall F1 score excluding this class, which is 82.4%.
@@
YOLOv8 [ultralyticsyolov8] 79.0 87.8 83.0 87.9